1)find three consecutive odd integers such that three times the first is 83 more than the sum of the second and the third
2)find four consecutive integers such that the sum of the last three integers is 20 more than two times the first

Answer by mananth(16946)

1)find three consecutive odd integers such that three times the first is 83 more than the sum of the second and the third
let the integers be x,x+2,x+4
3x= 83+x+2+x+4
3x=89+2x
3x-2x=89
x= 89
...
89,91,93
2)find four consecutive integers such that the sum of the last three integers is 20 more than two times the first
let the integers be x,x+1,x+2,x+3
...
x+1+x+2+x+3=2x+20
3x+6=2x+20
3x-2x=20-6
x=14
...
14,15,16,17
...
m.ananth@hotmail.ca